Can someone please help me solve this problem. Karen's average score for two test is 90. What score must Karen earn on a third test to raise her average to 92?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

(90+90+X)3=92
(180+X)/3=92
180+X=3*92
180+X=276
X=276-180
X=96 ANS. FOR THE THIRD TEST SCORE.
PROOF:
(90+90+96)/3=92
276/3=92
276=3*92
276=276

Answer by Alan3354(69443)